OpenStudy (anonymous):

find the domain of f(x)-x^2 over sqrt 25-x^2 and g(x)=x+5 over x-5

OpenStudy (anonymous):

f(x)= x^2 over sqrt 25-x^2 and g(x)=x+5 over x-5